OpenStudy (anonymous):

How do you do this? Find the side of a square whose diagonal is of the given measure. Given = 8 miles

OpenStudy (anonymous):

relation between side and diagonal of a square is : \[diagonal = \sqrt{side}\]

OpenStudy (anonymous):

sorry that is \sqrt{2} side

OpenStudy (anonymous):

\[side \sqrt{2}\]

OpenStudy (anonymous):

So would it just be